OpenStudy (anonymous):

Josie rode her bike 18 miles at 12 mi/h and then rode another 24 miles at 15 mi/h. How long did she ride in all? A. 3.1 hours B. 27 hours C. 15 hours D. 6 hours

OpenStudy (anonymous):

@Babynini @keshaun455 @viktorP554 @awkwardpanda @Afrodiddle @AngusV @AmyRoseRules

OpenStudy (afrodiddle):

Hello! :)

OpenStudy (anonymous):

Hi could you help me solve this problem?

OpenStudy (afrodiddle):

This is very simple; we are trying to find how many hours she rode in total today, so our formula is : Miles / miles per hour For example, if she rode 25 miles in the day, going at a constant speed of 12 we would divide 25 by 12 we would find out that she rode 2.1 hours. (2.08, but I estimated)

OpenStudy (afrodiddle):

That means what you need to do is divide the miles by the mph earlier that day, and then add it up with the later of the day. Miles / mph + Miles / mph = hours

OpenStudy (afrodiddle):

Do you get it? :)

OpenStudy (anonymous):

So i would divide 24 by?

OpenStudy (afrodiddle):

How many mph was she going when she rode 24? :)

OpenStudy (texaschic101):

time = distance / speed

OpenStudy (anonymous):

15

OpenStudy (afrodiddle):

So, so far we know that our equation is Miles / mph + 24 / 15 = hours

OpenStudy (afrodiddle):

What is the other Miles / mph ?

OpenStudy (anonymous):

12

OpenStudy (afrodiddle):

Miles / 12 What is the "miles" ?

OpenStudy (texaschic101):

t = d/t t = 18/12 + 24/15

OpenStudy (texaschic101):

oops...t = d/s

OpenStudy (afrodiddle):

Lol Times = distance / time :P

OpenStudy (texaschic101):

lol...way too many formulas in math to remember

OpenStudy (afrodiddle):

I never payed attention to formulas as a kid o.o I always just did it logically in my head.... I would always studder when teacher asked to write out my work because I did it all mentally lol.

OpenStudy (texaschic101):

well...as long as you got the right answer...I don't think it really matters

OpenStudy (afrodiddle):

Nope, they always yelled at me. o.o "BLAH BLAH BLAH" Like I copied off of someone else, even if I was the first one done. Lol... Elementary school logic.

OpenStudy (afrodiddle):

Ashley are you there?

OpenStudy (anonymous):

Yes

OpenStudy (afrodiddle):

Did you find the answer? :)